OpenStudy (anonymous):

What is the y-intercept of the line? y = (5/3)x - 5

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

y intercept is what you get when \(x=0\) and when \(x=0\) you get \(y=-5\)

OpenStudy (anonymous):

x=0 for y=(5/3)0-5 y=0-5 y=-5

OpenStudy (anonymous):

y=mx+b where b=y intercept